# Can LLM-Augmented autonomous agents cooperate?, An evaluation of their cooperative capabilities through Melting Pot

### v1

@@ Abstract
As the field of AI continues to evolve, a significant dimension of this progression is the development of Large Language Models and their potential to enhance multi-agent artificial intelligence systems. This paper explores the cooperative capabilities of Large Language Model-augmented Autonomous Agents (LAAs) using the well-known Meltin Pot environments along with reference models such as GPT4 and GPT3.5. Preliminary results suggest that while these agents demonstrate a propensity for cooperation, they still struggle with effective collaboration in given environments, emphasizing the need for more robust architectures. The study’s contributions include an abstraction layer to adapt Melting Pot game scenarios for LLMs, the implementation of a reusable architecture for LLM-mediated agent development – which includes short and long-term memories and different cognitive modules, and the evaluation of cooperation capabilities using a set of metrics tied to the Melting Pot’s ”Commons Harvest” game. The paper closes, by discussing the limitations of the current architectural framework and the potential of a new set of modules that fosters better cooperation among LAAs.
@@
The increased presence and relevance of AI agents within everyday spheres such as self-driving vehicles and customer service necessitates these entities being equipped with the appropriate capabilities to facilitate cooperation with humans and its AI counterparts. While noteworthy strides have been made in advancing individual intelligence components within AI agents, expanding the research focus to enhancing their social intelligence – the ability to effectively collaborate within group settings to solve prevalent problems – is now timely. This pivot aligns with the rapid progression of AI research presenting fresh prospects for fostering cooperation, drawing on insights from social choice theory and the development of social systems [1].
@@
Research into AI agents presents an avenue for generating intelligent technologies that embody more human-like features and are compatible with humans, a far cry from solipsistic approaches that overlook agent interactions. A promising illustration of this approach is the Melting Pot, an AI research tool designed to foster collaborative efforts within multi-agent artificial intelligence via canonical test scenarios. These environments emphasize non-trivial, learnable, and measurable cooperation by pairing a physical environment (a ”substrate”) with a reference set of co-players (a ”background population”) [2]. The environments fostered interdependence between the individuals involved.
@@
Research on AI agents has also recently been permeated by the leaps and bounds of Large Language Models (LLMs). The increasing success of LLMs encourages further exploration into LLM-augmented Autonomous Agents (LAAs). LAAs is an avenue of research that is still emerging, with limited explorations currently available [3, 4, 5, 6, 7, 8]. Something common in these works is that a clear need is established. To achieve success, LAAs have to rely on an architecture that can recall relevant events, reflect on such memories to generalize and draw a higher level of inferences, and utilize those reasonings to develop timely and long-term plans [9].
@@
These architectures offer specialized modules for specific tasks, utilizing meticulously crafted prompts and flows to perform complicated tasks and navigate intricate environments. Human behavior replication has been observed in some of these architectures, notably by Park et al. [9], who managed to create convincingly realistic human behavior in simulated environments. Similar frameworks utilized by Voyager [10] enabled an agent to navigate the Minecraft world and independently develop tools and skills. MetaGPT [11] introduced a framework allowing for the generation of fully functional programs, simulating a business-like environment with distinct roles and predefined agent interactions.
@@
Despite significant advancements in the field, the potential for cooperative abilities in (LAAs) has been somewhat neglected in current research. These capabilities could, however, be paramount in empowering these agents to perform innovative tasks and succeed in complex environments. This study represents an initial exploration into the inherent cooperative capabilities of LAAs. We employ an evaluation framework that includes a communication interface of scenarios from the Melting Pot project[2] (in which artificial agents co-exist in environments where social dilemmas can arise), the recent architecture proposed by Park et al.[9], as well as reference Large Language Models (LLMs) such as GPT4 and GPT3.5. Our results hint towards the capability for cooperative behavior, based on simple natural language definitions and cooperation metrics tailored to the chosen Melting Pot scenario. While the agents showed a propensity to cooperate, their actions did not demonstrate a clear understanding of effective collaboration within the given environment. Consequently, our analysis underscores the necessity for more robust architectures that can foster better collaboration in LAAs.
@@
Implementing a reusable architecture for the development of LAAs employing the modules proposed in [9]. This architecture includes short and long-term memories and cognitive modules of perception, planning, reflection, and action.
@@ 2 Related Work
A notable challenge with LLMs is their susceptibility to hallucinations and gaps in knowledge regarding recent events or specific subjects. Such constraints diminish their practicality, as they remain confined to the information acquired during training without the capability to assimilate new data. To address this issue, Shick et al. [7] introduced an early solution named Toolformer. This model was trained to discern when and how to invoke APIs (tools) to enhance the LLM’s performance across various tasks. The dataset was self-supervised, with API calls incorporated only when they positively impacted the model’s performance. This methodology empowered the model to determine the relevance and optimal execution of API calls.
@@
However, for executing more intricate tasks, merely invoking tools may be insufficient. Toolformer lacks the capability to reason about the rationale behind API calls and does not receive comprehensive environmental feedback to guide its subsequent actions toward achieving a goal. Recognizing this gap, the prompt-based paradigm ReAct [8], developed by Yao et al., integrates reasoning with action. By providing contextual prompt examples, ReAct guides the LLM on when to engage in reasoning and when to act, resulting in enhanced performance compared to approaches that employ reasoning or action in isolation. Furthermore, to enable the agent to learn from its errors, Shinn et al. [6] expanded the ReAct framework by incorporating a self-reflection module. This addition offers verbal feedback on past unsuccessful attempts, facilitating performance enhancement in subsequent trials.
@@
Conversely, drawing inspiration from the emulation of authentic human behavior, Park et al. [9] devised an intricate agent framework. This architecture boasts a cognitive sequence structured around modules primarily anchored by diverse prompts. Leveraging distinct prompts optimizes LLM performance, enabling specialized techniques for specific tasks. Notably, memory holds a pivotal position within this framework, preserving the agent’s experiences and insights. The ability to retrieve these memories diversely amplifies their utility across multiple objectives. Moreover, Wang et al. developed the Voyager architecture [10], enabling autonomous gameplay in Minecraft. This advanced framework empowers the agent to autonomously curate a discovery agenda. Remarkably, the architecture can also create its own APIs, write corresponding code, verify API functionality, and store it in a vector database for future utilization.
@@
More recently, efforts have been directed towards enhancing agent performance through multi-agent frameworks that utilize different instances of LLMs to independently perform roles or tasks. Du et al. [3] demonstrated this through a framework designed to engage different LLM instances in a debate, aiming to improve the factuality and accuracy of the responses. Subsequently, Hong et al. further capitalized on the potential of multiple agents. They allocated specific roles to each agent, accompanied by a sequence of predefined tasks with clear input and output expectations. These tasks establish a structured interaction pathway between agents, enabling them to achieve user-defined objectives. Demonstrating its efficacy in software-related tasks, this framework, inspired by the operational dynamics of conventional software companies, attained state-of-the-art performance in the HumanEval and MBPP benchmarks.
@@
Similarly, Liu et al. [4] introduced the BOLAA framework. This system orchestrates multi-agent activity by defining specialized agents overseen by a central controller. The controller’s role is pivotal: it selects the most suitable agent for a given task and facilitates communication with it. Additionally, Zhang et al. [5] delved into multi-agent architectures, exploring the influence of social traits and collaborative strategies on different datasets.
@@ 3.1 Experimental setup
Environment: This paper utilizes a scenario sourced from Melting Pot [2], a research tool developed by DeepMind for the purpose of experimentation and evaluation within the realm of multi-agent artificial intelligence. The scenarios within Melting Pot are specifically crafted to establish social situations in which the ability of the agents to solve conflict is challenged and characterized by significant interdependence among the involved agents.
@@
The LLM agents possess the capacity to execute high-level actions in each round. These actions include immobilize player (player_name) at (x, y), go to position (x, y), stay put, and explore (x, y). They enable the agents to zap other players, navigate to predefined positions on the map, stay in the same position, and explore the world respectively. On the contrary, bots, characterized as agents trained through reinforcement learning, perform one movement for every two movements made by any of the LLM agents. The policies governing the bots lead them to engage in unsustainable harvesting practices and instigate attacks against other agents in close proximity.
In general, maximizing the welfare population for this scenario would require the LLM agents to restrain themselves from eating the last apple in each of the apple trees, and to attack the bots or agents that take the apples in an unsustainable way to avoid the depletion of the apples.
Simulation: In a simulation, each episode of the game involves the participation of a predetermined quantity of LLM agents and bots. The LLM agents take a high-level action on their turn and proceed to execute it until all three LLM agents have completed their respective high-level actions. Meanwhile, the bots are in constant motion, executing a move for every two moves made by any of the agents (note that a high-level action typically comprises more than one movement). The simulation concludes either upon reaching a maximum predetermined number of rounds (100 typically) or prematurely if all the apples in the environment are consumed.
@@ 3.2 Adapting the environment to LLM agents
The Melting Pot scenarios consist of several two-dimensional layers accommodating various objects, each with its own custom logic. While initially, a matrix with distinct symbols seemed the most intuitive way to communicate the game state to the LLMs, it proved challenging for LLMs like GPT3.5 or GPT4 to interpret and reason about the spatial information provided by the position of objects in the matrix. To address this issue, we opted to develop an observation generator tailored to this particular environment. In this generator, every relevant object receives a natural language description, supplemented by coordinates expressed as a vector , denoting row and column respectively. Moreover, some relevant state changes are captured while an agent waits for its turn, and these changes are also captured and communicated to the agents. The complete list of descriptions generated for the objects and events of this environment is shown in Appendix A.
@@ 4 LLM agent architecture
The design of the LLM agents predominantly drew upon the Generative Agents architecture [9]. This choice was motivated by its comprehensive nature, positioning it as one of the most versatile architectures for agents that could be readily tailored to various tasks. While the Voyager architecture [10] also presented a viable option, its efficacy was somewhat limited due to its inherent inflexibility. Voyager constructs agent actions dynamically during gameplay, involving the generation and validation of code to execute actions in the environment. In the context of our specific case, it was deemed preferable to externalize actions from the architecture to enhance simplicity.
@@ 4.1 Memory structures
Long-Term Memory: This repository stores observations of the environment and the various thoughts generated by the agent in its cognitive modules. Leveraging the ChromaDB vector database, memories are stored, and the Ada OpenAI model generates contextual embeddings, enabling the agent to retrieve memories relevant to a given query.
Short-Term Memory: To facilitate rapid retrieval of specific memories or information, a Python dictionary is utilized. This dictionary stores information that must always be readily available to the agent, such as its name, as well as data that undergoes constant updates, such as current observations of the world.
Spatial Memory: Given the agent’s navigation requirements in a grid world environment, spatial information becomes pivotal. This includes the agent’s position and orientation. To support effective navigation from one point to another, utility functions are implemented to aid the agent in spatial awareness and movement.
@@ 4.2 Cognitive modules
Perception module: The initial stage in the cognitive sequence is the Perception Module. This module is tasked with assimilating raw observations from the environment. These observations serve as a comprehensive snapshot of the current state of the world, offering insights into the items within the agent’s observable window.
@@
Ultimately, the Perceive module determines whether an agent should initiate a response based on the current observations. During this stage, the agent assesses its existing plan and queued actions to ascertain their suitability. It evaluates whether it is appropriate to proceed with the current course of action or if the observed conditions warrant the development of a new plan and the generation of corresponding actions for execution. The complete prompt is shown in Appendix C.
Planning module: This module comes into play once observations have been sorted and filtered. The Planning modules leverage the amalgamation of current observations, the existing plan, the contextual understanding of the world, reflections from the past, and rationale to meticulously craft a newly devised plan. This plan intricately outlines the high-level behavior expected from the agent and delineates the goals the agent will diligently pursue. For the complete prompt refer to Appendix D.
Reflection module: The Reflect module is designed to facilitate profound contemplation on observations and fellow agent thoughts at a higher cognitive level. Activation of this module is contingent upon reaching a predetermined threshold of accumulated observations. In our experimental setup, reflections were initiated every 30 perceived observations, roughly translating to three rounds in the game. The Reflect model comprises two key stages:

### v2

@@ Abstract
As the field of AI continues to evolve, a significant dimension of this progression is the development of Large Language Models (LLMs) and their potential to enhance multi-agent artificial intelligence systems. This paper explores the cooperative capabilities of Large Language Model-augmented Autonomous Agents (LAAs) using the well-known Melting Pot environments along with reference models such as GPT-4 and GPT-3.5. Preliminary results suggest that while these agents demonstrate a propensity for cooperation, they still struggle with effective collaboration in given environments, emphasizing the need for more robust architectures. The study’s contributions include an abstraction layer to adapt Melting Pot game scenarios for LLMs, the implementation of a reusable architecture for LLM-mediated agent development (which includes short and long-term memories and different cognitive modules), and the evaluation of cooperation capabilities using a set of metrics tied to the Melting Pot’s “Commons Harvest” game. The paper closes by discussing the limitations of the current architectural framework and the potential of a new set of modules that fosters better cooperation among LAAs.
@@ 1 Introduction
The increased presence and relevance of AI agents within everyday spheres such as self-driving vehicles and customer service necessitates, these entities being equipped with the appropriate capabilities to facilitate cooperation with humans and its AI counterparts. While noteworthy strides have been made in advancing individual intelligence components within AI agents, expanding the research focus to enhance their social intelligence —the ability to effectively collaborate within group settings to solve prevalent problems–- is now timely. This pivot aligns with the rapid progression of AI research presenting fresh prospects for fostering cooperation, drawing on insights from social choice theory and the development of social systems (Dafoe et al., 2020).
Moreover, using AI to manage open innovation processes provides a framework for improving AI’s social intelligence. By integrating AI to perform key functions such as mapping the innovation landscape, coordinating diverse knowledge inputs, and ensuring collaboration aligns with collective objectives, we can better understand and leverage AI’s capabilities in facilitating complex collaborative efforts within and across communities (Broekhuizen et al., 2023).
Cooperation requires the presence of two or more agents who, based on mutual understanding, engage in collaborative actions. From the perspective of evolution, cooperation has played a crucial role in the survival of species (Pennisi, 2009; Dale et al., 2020). It has facilitated the development of social structures that influence our surroundings and has provided solutions to complex issues, such as social dilemmas (Gross et al., 2023). Unveiling the processes that enable the evolution of cooperative behavior in communities is regarded as an important challenge that scientists should tackle (Pennisi, 2009).
Mathematical and computational models of cooperation have been proposed to study the mechanisms that facilitate the emergence of cooperation among artificial agents engaged in collaborative actions with the aim of improving their joint welfare (Dafoe et al., 2020). For example, matrix games have provided a useful tool to conduct research on social dilemmas for years (Axelrod and Hamilton, 1981), in which the decision to either cooperate or defect is restricted to an atomic action. Additionally, multi-agent reinforcement learning has been used to study the process of learning cooperative policies in complex scenarios where temporally extended social dilemmas arise (Leibo et al., 2017, 2021; McKee et al., 2023; Rios et al., 2023).
@@
Research into AI agents presents an avenue for generating intelligent technologies that embody more human-like features and are compatible with humans, a far cry from solipsistic approaches that overlook agent interactions. A promising illustration of this approach is the Melting Pot, an AI research tool designed to foster collaborative efforts within multi-agent artificial intelligence via canonical test scenarios. These environments emphasize non-trivial, learnable, and measurable cooperation by pairing a physical environment (a “substrate”) with a reference set of co-players (a “background population”) (Agapiou et al., 2023). The environments fostered interdependence between the individuals involved.
@@
Research on AI agents has also recently been permeated by the leaps and bounds of Large Language Models (LLMs). The increasing success of LLMs encourages further exploration into LLM-augmented Autonomous Agents (LAAs). LAAs represent an avenue of research that is still emerging, with limited explorations currently available (Du et al., 2023; Liu et al., 2023; Zhang et al., 2023; Shinn et al., 2023; Schick et al., 2023; Yao et al., 2023). Something common in these works is that a clear need is established. To achieve success, LAAs have to rely on an architecture that can recall relevant events, reflect on such memories to generalize and draw a higher level of inferences, and utilize those reasonings to develop timely and long-term plans (Park et al., 2023).
@@
These architectures offer specialized modules for specific tasks, utilizing meticulously crafted prompts and flows to perform complicated tasks and navigate intricate environments. Human behavior replication has been observed in some of these architectures, notably by Park et al. (2023), who managed to create convincingly realistic human behavior in simulated environments. Similar frameworks utilized by Voyager (Wang et al., 2023) enabled an agent to navigate the Minecraft world and independently develop tools and skills. MetaGPT (Hong et al., 2023) introduced a framework allowing for the generation of fully functional programs, simulating a business-like environment with distinct roles and predefined agent interactions.
@@
Despite significant advancements in the field, the potential for cooperative abilities in Large Language Models augmented Autonomous Agents (LAAs) has been somewhat neglected in current research. These capabilities could, however, be paramount in empowering these agents to perform innovative tasks and succeed in complex environments. This study represents an initial exploration into the inherent cooperative capabilities of LAAs. We employ an evaluation framework that includes a communication interface of scenarios from the Melting Pot project (Agapiou et al., 2023) (in which artificial agents co-exist in environments where social dilemmas can arise), the recent architecture proposed by Park et al. (2023), as well as reference Large Language Models (LLMs) such as GPT-4 and GPT-3.5. Our results hint towards the capability for cooperative behavior, based on simple natural language definitions and cooperation metrics tailored to the chosen Melting Pot scenario. While the agents showed a propensity to cooperate, their actions did not demonstrate a clear understanding of effective collaboration within the given environment. Consequently, our analysis underscores the necessity for more robust architectures that can foster better collaboration in LAAs.
@@
Implementing a reusable architecture for the development of LAAs employing the modules proposed in Generative Agents (Park et al., 2023). This architecture includes short- and long-term memories and cognitive modules of perception, planning, reflection, and action. Our project can be found at https://github.com/Cooperative-IA/CooperativeGPT
@@ 2 Related Work
A notable challenge with LLMs is their susceptibility to hallucinations and gaps in knowledge regarding recent events or specific subjects. Such constraints diminish their practicality, as they remain confined to the information acquired during training without the capability to assimilate new data. To address this issue, Schick et al. (2023) introduced an early solution named Toolformer. This model was trained to discern when and how to invoke APIs (tools) to enhance the LLM’s performance across various tasks. The dataset was self-supervised, with API calls incorporated only when they positively impacted the model’s performance. This methodology empowered the model to determine the relevance and optimal execution of API calls.
@@
However, for executing more intricate tasks, merely invoking tools may be insufficient. Toolformer lacks the capability to reason about the rationale behind API calls and does not receive comprehensive environmental feedback to guide its subsequent actions toward achieving a goal. Recognizing this gap, the prompt-based paradigm ReAct, developed by Yao et al. (2023), integrates reasoning with action. By providing contextual prompt examples, ReAct guides the LLM on when to engage in reasoning and when to act, resulting in enhanced performance compared to approaches that employ reasoning or action in isolation. Furthermore, to enable the agent to learn from its errors, Shinn et al. (2023) expanded the ReAct framework by incorporating a self-reflection module. This addition offers verbal feedback on past unsuccessful attempts, facilitating performance enhancement in subsequent trials.
@@
Conversely, drawing inspiration from the emulation of authentic human behavior, Park et al. (2023) devised an intricate agent framework. This architecture boasts a cognitive sequence structured around modules primarily anchored by diverse prompts. Leveraging distinct prompts optimizes LLM performance, enabling specialized techniques for specific tasks. Notably, memory holds a pivotal position within this framework, preserving the agent’s experiences and insights. The ability to retrieve these memories diversely amplifies their utility across multiple objectives. Moreover, Wang et al. (2023) developed the Voyager architecture, enabling autonomous gameplay in Minecraft. This advanced framework empowers the agent to autonomously curate a discovery agenda. Remarkably, the architecture can also create its own APIs, write corresponding code, verify API functionality, and store it in a vector database for future utilization.
@@
More recently, efforts have been directed towards enhancing agent performance through multi-agent frameworks that utilize different instances of LLMs to independently perform roles or tasks. Du et al. (2023) demonstrated this through a framework designed to engage different LLM instances in a debate, aiming to improve the factuality and accuracy of the responses. Subsequently, Hong et al. further capitalized on the potential of multiple agents. They allocated specific roles to each agent, accompanied by a sequence of predefined tasks with clear input and output expectations. These tasks establish a structured interaction pathway between agents, enabling them to achieve user-defined objectives. Demonstrating its efficacy in software-related tasks, this framework, inspired by the operational dynamics of conventional software companies, attained state-of-the-art performance in the HumanEval and MBPP benchmarks.
@@
Similarly, Liu et al. (2023) introduced the BOLAA framework. This system orchestrates multi-agent activity by defining specialized agents overseen by a central controller. The controller’s role is pivotal: it selects the most suitable agent for a given task and facilitates communication with it. Additionally, Zhang et al. (2023) delved into multi-agent architectures, exploring the influence of social traits and collaborative strategies on different datasets. Likewise, Ni and Buehler (2024) developed “MechAgents,” a system employing multiple dynamically interacting LLMs to autonomously solve mechanics tasks. This framework not only shows that self-correcting and mutually correcting code can work between AI agents, but it also makes it easier for the agents to combine physics-based modeling with domain-specific knowledge. This opens up a new way to automate and improve the way engineers solve problems.
@@ 3.1 Experimental setup
3.1.1 Environment
This paper utilizes a scenario sourced from Melting Pot (Agapiou et al., 2023), a research tool developed by DeepMind for the purpose of experimentation and evaluation within the realm of multi-agent artificial intelligence. The scenarios within Melting Pot are specifically crafted to establish social situations in which the ability of the agents to solve conflict is challenged and characterized by significant interdependence among the involved agents.
@@ 3.1.1 Environment
The LLM agents possess the capacity to execute high-level actions in each round. These actions include: immobilize player (player_name) at (x, y), go to position (x, y), stay put, and explore (x, y). They enable the agents to zap other players, navigate to predefined positions on the map, stay in the same position, and explore the world, respectively. On the contrary, bots, characterized as agents trained through reinforcement learning, perform one movement for every two movements made by any of the LLM agents. The policies governing the bots lead them to engage in unsustainable harvesting practices and instigate attacks against other agents in close proximity.
In general, maximizing the welfare of the population for this scenario would require the LLM agents to restrain themselves from eating the last apple on each of the apple trees, and to attack the bots or agents that harvest the apples in an unsustainable way to avoid the depletion of the apples.
3.1.2 Simulation
In a simulation, each episode of the game involves the participation of a predetermined quantity of LLM agents and bots. The LLM agents take a high-level action on their turn and proceed to execute it until all three LLM agents have completed their respective high-level actions. Meanwhile, the bots are in constant motion, executing a move for every two moves made by any of the agents (note that a high-level action typically comprises more than one movement). The simulation concludes either upon reaching a maximum predetermined number of rounds (typically 100) or prematurely if all the apples in the environment are consumed.
@@ 3.2 Adapting the environment to LLM agents
The Melting Pot scenarios consist of several two-dimensional layers accommodating various objects, each with its own custom logic. While initially, a matrix with distinct symbols seemed the most intuitive way to communicate the game state to the LLMs, it proved challenging for LLMs like GPT-3.5 or GPT-4 to interpret and reason about the spatial information provided by the position of objects in the matrix. To address this issue, we opted to develop an observation generator tailored to this particular environment. In this generator, every relevant object receives a natural language description, supplemented by coordinates expressed as a vector , denoting row and column respectively. Moreover, some relevant state changes are captured while an agent waits for its turn, and these changes are also captured and communicated to the agents. The complete list of descriptions generated for the objects and events of this environment is shown in Appendix A.
@@ 4 LLM agent architecture
The design of the LLM agents predominantly drew upon the Generative Agents architecture (Park et al., 2023). This choice was motivated by its comprehensive nature, positioning it as one of the most versatile architectures for agents that could be readily tailored to various tasks. While the Voyager architecture (Wang et al., 2023) also presented a viable option, its efficacy was somewhat limited due to its inherent inflexibility. Voyager constructs agent actions dynamically during gameplay, involving the generation and validation of code to execute actions in the environment. In the context of our specific case, it was deemed preferable to externalize actions from the architecture to enhance simplicity.
@@ 4.1 Memory structures
4.1.1 Long-Term Memory
This repository stores observations of the environment and various thoughts generated by the agent in its cognitive modules. Leveraging the ChromaDB vector database, memories are stored and the Ada OpenAI model generates contextual embeddings, enabling the agent to retrieve memories relevant to a given query.
4.1.2 Short-Term Memory
To facilitate rapid retrieval of specific memories or information, a Python dictionary is utilized. This dictionary stores information that must always be readily available to the agent, such as its name, as well as data that undergoes constant updates, such as current observations of the world.
4.1.3 Spatial Memory
Given the agent’s navigation requirements in a grid world environment, spatial information becomes pivotal. This includes the agent’s position and orientation. To support effective navigation from one point to another, utility functions are implemented to aid the agent in spatial awareness and movement.
@@ 4.2 Cognitive modules
4.2.1 Perception module
The initial stage in the cognitive sequence is the Perception Module. This module is tasked with assimilating raw observations from the environment. These observations serve as a comprehensive snapshot of the current state of the world, offering insights into the items within the agent’s observable window.
@@ 4.2.1 Perception module
Ultimately, the Perceive module determines whether an agent should initiate a response based on the current observations. During this stage, the agent assesses its existing plan and queued actions to ascertain their suitability. It evaluates whether it is appropriate to proceed with the current course of action or if the observed conditions warrant the development of a new plan and the generation of corresponding actions for execution. The complete prompt is shown in C.
4.2.2 Planning module
This module comes into play once observations have been sorted and filtered. The Planning module leverages the amalgamation of current observations, the existing plan, the contextual understanding of the world, reflections from the past, and rationale to meticulously craft a newly devised plan. This plan intricately outlines the high-level behavior expected from the agent and delineates the goals the agent will diligently pursue. For the complete prompt, refer to D.
4.2.3 Reflection module
The Reflection module is designed to facilitate profound contemplation on observations and thoughts from fellow agents at a higher cognitive level. Activation of this module is contingent upon reaching a predetermined threshold of accumulated observations. In our experimental setup, reflections were initiated after every 30 perceived observations, roughly translating to three rounds in the game. The Reflection module comprises two key stages: